1.    Two sides of a square are on the line x + 10y – 10 = 0 and x + 10y – 5 = 0. Find the area of the square.



1
Expert's answer
2021-11-06T01:36:41-0400

The distance between two parallel lines "ax+by+c=0" and "ax+by+d=0"


"l=\frac{|c-d|}{\sqrt{a^2+b^2}}"


So, we have "l=\frac{|c-d|}{\sqrt{a^2+b^2}}=\frac{|-10-(-5)|}{\sqrt{1^2+10^2}}=0.4975"


The area of the square "A=0.4975^2=0.2475" . Answer